Homework Help-Physics

What is the electric potential a distance of 2.6 t
Answer #1

A,. Electric potentail V = Kq/r

so V = 9e9 *1.6*10^-19/2.6*10^-15

V = 5.53*10^5 Volts



b. potentail enrnergy U = Vq

U = 5.53*10^5 * 1.6*10^-19

U = 8.848 *10^-14 J
Answer #3

(a) V = kq/r

=> V = 8.98 * 10^9 * 1.67 * 10^(-19) /(2.6 * 10^(-15))

=> V = 5.7679 * 10^5 J/C

(b) U = Kq1q2/r

=> U = 8.98 * 10^9 * 1.67 * 10^(-19)* 1.67 * 10^(-19)/(2.6 * 10^(-15))

=> U = 9.63 * 10^(-14) J